To buy a computer system, a customer can choose one of 4 monitors, one of 2 keyboards, one of 4 computers and one of 3 printers. Determine the number of possible systems that a customer can choose from.
1
Expert's answer
2020-05-26T18:57:43-0400

If we have to select r items from a total of n items then the no. of ways for doing that is given by "^nC_r"


No. of options the customer has to choose from

"^4C_1\times^2C_1\times^4C_1\times^3C_1=4\times2\times4\times3=96"